# Wiedemann-Franz law violation domain for graphene and nonrelativistic systems

### v2

@@ II.2 An approach towards the fluid description
The previous subsection provides the expression of $\kappa$ and $\sigma$ using a standard solid state physics framework, where no fluid concept has been entered. So, we can call that description a non-fluid description. In the present section, we will try to build an approach towards a fluid description by using relaxation time approximation-based Boltzmann transport equations. Here, we are using the terminology ”towards the fluid description” because the actual phenomenological fluid description [21] for explaining the Wiedemann-Franz law violation in the experimental 2D graphene system is quite rigorous. Aim of the present work is to zoom in the intermediate journey from non-fluid to fluid description. Interestingly, we will get a Wiedemann-Franz law violating domain in $T-\epsilon_{F}$ plane for all cases. Let us first address the brief calculations of the Lorenz ratio for 2D-G systems; then, we will write down its final expressions for other possible systems like 2D-NR, 3D-NR, and 3D-G instead of repeating similar calculations.
@@
It is relaxation time approximation (RTA) based on Boltzmann’s transport equation (BTE) [42]
@@
Here, we consider the approximation $[C_{V}]_{e}\approx\frac{\partial\epsilon}{\partial T}$, and one can expect again two possible definitions of specific heat as discussed in the earlier section. From Eq. (44), we can get the form of $\delta f_{\sigma}$ and $\delta f_{\kappa}$ as
@@
$\frac{L_{G2}^{2D}}{L_{0}}=\frac{12}{\pi^{2}}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}$ (52)
@@
2D-NR After doing the same calculation as the previous one for a 2-dimensional nonrelativistic system, the final general expressions of the electrical and two possible thermal conductivities and their corresponding Lorenz ratios $L_{NR}^{2D}/L_{0}$ are given by as below:
$f_{\frac{5}{2}}\left(A\right)T^{\frac{5}{2}},$ (64)
$f_{\frac{5}{2}}\left(A\right)T^{\frac{5}{2}},$ (65)
@@ III Results
Figure 1: The Lorenz ratio regarding the chemical potential for non-fluid descriptions of graphene case (a) for 3D and (b) for 2D expressions.
@@
Figure 2: The Lorenz ratio regarding the chemical potential for non-fluid descriptions of nonrelativistic case (a) for 3D and (b) for 2D expressions.
@@
In the above section, we have calculated all the general expressions of electrical and thermal conductivity and their ratios to check the validity of the Wiedemann-Franz law for all possible systems like 2D and 3D graphene (G) and nonrelativistic (NR) cases. Using those final expressions, the present section is intended to go with their numerical estimations. During the result generation, we will use natural unit $k_{B}=\hbar=c=1$ for our convenience. Dimensionless quantities Lorenz ratio ($L/L_{0}$) and $\epsilon_{F}/T$ will be taken as the y and x axes of all graphs.
@@
Figure 3: Left: Lorenz ratio of 2D-G systems by using fluid-type expression. Right: Comparison of specific heat and fluid aspect in 2D-G system.
@@
Figure 4: Left: The Lorenz ratio of fluid and non-fluid for 2D-G system. Right: The Lorenz ratio of the mixture of fluid and non-fluid with fitting parameters for 2D-G system.
@@
In Fig. 1 and Fig. 2, we explored the Lorenz ratio ($\frac{L}{L_{0}}$) by using non-fluid (NF) type expressions. Fig. 1(a) and (b) have shown the Lorenz ratio versus $\frac{\epsilon_{F}}{T}$ for 3D-G and 2D-G systems respectively. By using the non-fluid (NF) type of Eq. (20) and its Sommerfeld limit Eq. (24), we plotted the black solid and dotted lines in Fig. 1(a). We used the specific heat ${[C_{V}]_{e1}}=\frac{\partial}{\partial T}\Big{(}\frac{U}{N}\Big{)}$, which is traditionally used in solid state or non-fluid systems. We are getting $L/L_{0}=1$ in the limit of $\epsilon_{F}/T\gg 1$ or Fermi Liquid (FL) regime of graphene case, which is as expected for most of the metal cases. The deviation of the Lorenz ratio from 1 has been started below $\frac{\epsilon_{F}}{T}\approx 10$ as we notice from the black solid line. We have also plotted the red solid line by using Eq. (21), where another definition of specific heat ${[C_{V}]_{e2}}$ is used in the thermal conductivity expression. In the case of SL of 3D-G (the red dotted line), the Lorenz ratio is three instead of one. For the general case (the red solid line), the Lorenz ratio deviates from this constant line $3$ below $\frac{\epsilon_{F}}{T}\approx 10$. For the 2D-G system, similar kinds of black and red solid lines are plotted in Fig. 1(b) by using Eq. (36) and Eq. (37). Their respective SLs by using Eq. (38) and Eq. (39), are plotted as black and red horizontal dotted lines. We considered the surface area $(S)$ in place of volume $(V)$ in the 2D system. The Lorenz ratio when we take ${[C_{V}]_{e2}}$, the ratio becomes 2 (the red-dotted line).
@@
The Fig. 2(a) and (b), similar to Fig. 1, display the Lorenz ratio in terms of $\frac{\epsilon_{F}}{T}$ for non-fluid expressions of NR case. Though the real metals have a constant chemical potential within the range, $\epsilon_{F}=2-10$ eV, we imagine that it can be tunable from 00 to $\infty$. In Fig. 2(a), we notice that the WF law is valid in the SL of the 3D metal case, expressed by Eq. (10) (green dot-dashed line). We have taken $T=60^{\circ}$ K ($\approx 0.005$ eV), so metal range $\epsilon_{F}=2$-$10$ eV will give the ratio $\frac{\epsilon_{F}}{T}=400$-$2000$ ($\gg 1$), which is quite good domain for SL. Therefore, being situated in the SL domain, metal always follows the WF law. A slight larger value $L=1.5L_{0}$ comes from Eq. (14) (blue dotted line), when specific heat ${[C_{V}]_{e2}}$ is taken. A similar kind of behavior can be seen in Fig. 2(b) for 2D-NR systems, but interestingly, SLs in both cases, given by Eqs. (31) and (32) are coincided at $L=L_{0}$ (blue-dotted and green dot-dashed lines). It probably indicates that if one builds 2D metal systems, then it will also follow the WF law as their Fermi energy range is again expected to be situated in the SL domain.
@@
Before going to the next graphical results, let us briefly analyze what we have learned from Fig. 1(a), (b) and Fig. 2(a), (b). Grossly, the non-fluid type expressions of Lorenz ratios for all cases - 3D-NR, 2D-NR, 3D-G, and 2D-G are saturated near 1 in the SL range, which means that the WF law is obeyed in the SL range. Among all four cases, 3D-NR and 2D-G are realistic cases. The former is applicable to metal systems, whose Fermi energies remain constant and within the SL range, so they always follow the WF law. Latter case - the non-fluid expressions of Lorenz ratio for 2D-G can be applicable for realistic graphene system with high doping concentration ($\epsilon_{F}$ is quite large or $\epsilon_{F}/T\gg 1$), where the WF law is obeyed well again [15]. However, by decreasing the doping concentration, when the graphene system approaches the charge-neutral points ($\epsilon_{F}$ is quite small or $\epsilon_{F}/T\ll 1$), the WF law violation domain is observed [15]. According to our non-fluid expressions, the WF law violation is also expected in the $\epsilon_{F}/T\ll 1$ domain but the outcome $L/L_{0}<1$ is not matching with experimental outcome $L/L_{0}>1$ by Crossno et al. [15]. A possible explanation of $L/L_{0}>1$ in the appearance of electron hydrodynamics [21] property in the domain $\epsilon_{F}/T<1$, which is popularly called the Dirac fluid domain. Present work is firstly intended to highlight the overlooked possibility of Wiedemann-Franz law violation of any systems like NR, G or UR even we consider the non-fluid expression of $\kappa$ and $\sigma$. Then, its next attempt is to show the transition from non-fluid (NF) to fluid (F) pictures in a systematic way. An approach towards the fluid description is addressed in formalism part Sec: (II.2). Let us come to the numerical estimation for fluid-type formalism.
@@
For numerical curves of fluid-type expressions, let us focus only on the 2D-G case. Although formalism of other cases with an Appendix is provided in the present article, one may go for result generation for graphene case. Using Eqs. (51) and (52), black and red solid lines are plotted in the left panel of Fig. 3. They are saturated at the values of 2 and 4, respectively, in the high doping or $\epsilon_{F}/T\gg 1$ domain. It means that if the fluid type expression is valid in this high doping domain 2D-G system, then the WF law may not be obeyed. However, the experimental data [15] claims that the WF law is well obeyed in the high doping or $\epsilon_{F}/T\gg 1$ domain. This means that the non-fluid expression of the Lorenz ratio should be applicable to this domain instead of fluid-type expression. The SL of Eqs. (51) and (52) will also be
@@
which are plotted by black and red dotted lines, respectively, in the left panel of Fig. 3. They blow up from their saturated values (2 and 4) when $\epsilon_{F}/T$ or doping concentration decreases. The experimental data of Crossno et al. [15] is pasted, where S1 and S2 represent the samples of the 2D-G system in terms of lower to higher values of minimum charge carrier density. Theoretically, zero charge carrier density can be reached at $\epsilon_{F}/T\rightarrow 0$ but experimentally, a minimum charge carrier density [15] will be achieved. So, from Fig.3(a), we notice that our fluid-based expression of $\frac{L_{G2}^{2D}}{L_{0}}$, given in Eq.(51), (52), can’t able to explain the experimental data points. It indicates that our formalism needs next level modification. Due to this step by step development of our formalism from (standard) non-fluid to (new directional) fluid description, we put the name of Sec: (II.2) as ”towards the fluid description”. Now if we revisit our RTA based kinetic theory formalism, then we can find that we have systematically considered the local thermalization concept of fluid. In this concept, we consider fluid as a large number fluid elements, having velocity $u^{\mu}$, temperature T and chemical potential $\epsilon_{F}$. Now, fluid elements will be at different position and time $(x^{\mu})$, so, considering $u^{\mu}(x^{\mu})$, $T(x^{\mu})$, $\epsilon_{F}(x^{\mu})$ as function of $x^{\mu}$ can map mathematically the local thermalization concept. Since present work is only focus on thermal and electrical conductivity, so only $T(x^{\mu})$ is considered and others $u^{\mu}(x^{\mu})$ and $\epsilon_{F}(x^{\mu})$ are ignored. Reader can see this consideration in Eq.(44), where the information of temperature gradient $\frac{\partial T}{\partial x^{i}}$ is entered. Now the another term specific heat $[C_{V}]_{e}$ $\approx\frac{\partial u}{\partial T}$, which is basically considered as temperature derivative of average energy of electron. This average energy of electron information can be modified into enthalpy per electron $h=\frac{u+P}{n}$ in fluid picture. Another point is that electron-hole plasma picture will be important when we go $\frac{\epsilon_{F}}{T}<1$ domain (towards charge neutrality point). So, we should take enthalpy per net electron $h=\frac{u+P}{n}$, where $u=u_{e}+u_{h}$, $P=P_{e}+P_{h}$ and $n=n_{e}-n_{h}$ will be respectively total internal energy density, pressure and net charge density. We have adopted the replacement
$[C_{V}]_{e}\approx\frac{\partial u}{\partial T}\,\rightarrow\,\frac{h}{T}=\frac{u+P}{nT}$ (70)
@@
$\left(\frac{L_{G1}^{2D}}{L_{0}}\right)=\frac{6}{\pi^{2}}\left(\frac{u+P}{nT}\right)\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}~{}.$ (72)
@@
  • Fluid (F) domain: Fluid based Lorenz ratio, given in Eq.(72) valid well with experimental data in the range $0.70.2<\frac{\epsilon_{F}}{T}<$ 0.7.
  • Non-Fluid (NF) domain: Non-Fluid based Lorenz ratio, given in Eq.(36), valid well with experimental data in the ranges (a) $102.5<\frac{\epsilon_{F}}{T}<$ 10 and (b) $10<\frac{\epsilon_{F}}{T}<\infty$. We can called subrange (b) as ohmic non-fluid or NF(Ohmic) domain, since we get WF law or $\frac{L}{L_{0}}=1$ for this $\frac{\epsilon_{F}}{T}>10$ domain, where Fermi-Dirac distribution function can be assumed as step function. However, for subrange (a), we get $\frac{L}{L_{0}}<1$ because assumption of step function is not further valid here. Corresponding domains are also marked in Fig.3(b), where we can notice that $C_{V}^{e}\propto\frac{T}{\epsilon_{F}}$ follows in NF(Ohmic) domain (due to validation of electron step function like distribution) but a transition from $C_{V}^{e}\propto\frac{T}{\epsilon_{F}}$ to $C_{V}^{e}=2k_{B}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg{]}$ is observed in NF domain ($102.5<\frac{\epsilon_{F}}{T}<$ 10). Reader can understand this transition from Eq.(34).
  • Mixed or (F+NF) domain: The domain $2.50.7<\frac{\epsilon_{F}}{T}<$ 2.5 is marked as mixed or (F+NF) domain as we have to go with replacement $C_{V}^{e}\rightarrow\frac{u+P}{nT}$ or transit from NF based quantity $C_{V}^{e}$ to fluid-based quantity $\frac{u+P}{nT}$. This domain may be considered as a grey zone and we may put an open question for further research- How this transition happens? Interestingly, reader can notice a $\frac{\epsilon_{F}}{T}$ independent trend approximately of $C_{V}^{e}$ and $\frac{u+P}{nT}$ within this (F+NF) domain.
@@
$\left(\frac{L}{L_{0}}\right)_{F/NF}=\frac{6}{\pi^{2}}\,\big{[}C_{V}^{e}\big{]}_{F/NF}\,\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\,(A=e^{x})~{},$ (76)
@@
Using Eq.(76), we have plotted blue and red solid curves to fit experimental data point $S_{1}$(blue circles) and $S_{1}$(red circles). For blue line, tuning parameters are taken as $x_{F}=0.2$, $x_{NF}=0.8$, $b_{F}=2$, $b_{NF}=0.7$ and for red line, tuning parameters are taken as $x_{F}=0.9$, $x_{NF}=2$, $b_{F}=2$, $b_{NF}=0.7$. By using a better version of x-dependent switching functions, one may go with better fitting of experimental data. However, instead of quantitative exact fitting, the core aim of the present work may be considered as its qualitative message or understanding on the F and NF domain and their transition.
@@
This qualitative message of transition from NF to F is in agreement with earlier theoretical works by Lucas et al. [21]. Though final expressions of Lorenz ratio of Ref. [21] and the present work are little different but both works have similar qualitative message- NF to F transition is associated with the WF law violation and enhancement of Lorenz ratio more than one. There are other works also which provide little different reasons for getting WF law violations. Ref. [20] has linked this fact with presence of band gap and bipolar diffusion, while Ref. [19] has built thermodynamical expression of electron-hole plasma in terms of polylogarithm function, which can provide a greater than one values of Lorenz ratio. So, many more theoretical research on this topic are expected in future for getting some converging understanding.
IV Summary and Interpretation

### v3

@@ II.2 An approach towards the fluid description
The previous subsection provides the expression of $\kappa$ and $\sigma$ using a standard solid state physics framework, where no fluid concept has been entered. So, we can call that the non-fluid description. In the present section, we will try to build an approach towards a fluid description by using relaxation time approximation-based Boltzmann’s transport equations. Here, we are using the terminology ”towards the fluid description” because the actual phenomenological fluid description [Theory_G_WF] for explaining the Wiedemann-Franz law violation in the experimental 2D graphene system is quite rigorous. Aim of the present work is to zoom in the intermediate journey from non-fluid to fluid description. Interestingly, we will get a Wiedemann-Franz law violating domain in $T-\epsilon_{F}$ plane for all cases. Let us first address the brief calculations of the Lorenz ratio for 2D-G systems; then, we will write down its final expressions for other possible systems like 2D-NR, 3D-NR, and 3D-G instead of repeating similar calculations (see Appendix E).
@@
It is relaxation time approximation (RTA) based on Boltzmann’s transport equation (BTE) [BTE]
@@
Here, we consider the approximation $C_{V}]_{e}\approx\frac{\partial\epsilon}{\partial T}$, and one can expect again two possible definitions of specific heat as discussed in the earlier section. From Eq. (20), we can get the form of $\delta f_{\sigma}$ and $\delta f_{\kappa}$ as
@@
$\delta f_{\kappa}=v^{i}\left(-\frac{\partial f_{0}}{\partial\epsilon}\right)[C_{V}]\left(\frac{\partial T}{\partial x^{i}}\right)\tau_{c}$ (22)
@@
$\kappa_{F/NF}^{G}=\frac{\tau_{c}}{\pi}[C_{V}]_{F/NF}\,f_{2}\left(A\right)T^{2}.$ (26)
$\kappa_{F/NF}^{Q}=\frac{24\ \tau_{c}}{\pi^{2}}[C_{V}]_{F/NF}\,f_{3}\left(A\right)T^{3}.$ (27)
@@
$\frac{L_{G2}^{2D}}{L_{0}}=\frac{12}{\pi^{2}}\Bigg[3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg]\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}$ (31)
@@ III Results
Figure 1: The Lorenz ratio regarding the chemical potential for non-fluid descriptions of graphene case (a) for 3D and (b) for 2D expressions
@@
Figure 2: The Lorenz ratio regarding the chemical potential for non-fluid descriptions of nonrelativistic case (a) for 3D and (b) for 2D expressions
@@
In the above section, we have calculated all the general expressions of electrical and thermal conductivity and their ratios to check the validity of the Wiedemann-Franz law fore 2D graphene (G) system. Using those final expressions, the present section is intended to go with their numerical estimations. During the result generation, we used the natural unit $k_{B}=\hbar=c=1$ for our convenience. Dimensionless quantities Lorenz ratio ($L/L_{0}$) and $\epsilon_{F}/T$ will be taken as the y and x axes of all graphs.
@@
Figure 3: Lorenz ratio of 2D-G system by using fluid-type expression versus $\epsilon_{F}/T$ (LEFT) and comparison of specific heat and fluid aspect in 2D-G system versus $\epsilon_{F}/T$ (RIGHT)
@@
Figure 4: Lorenz ratio of fluid and non-fluid for 2D-G system versus $\epsilon_{F}/T$ (LEFT) and Lorenz ratio of the mixture of fluid and non-fluid with fitting parameters for 2D-G system versus $\epsilon_{F}/T$ (RIGHT)
@@
In Fig. 1 and Fig. 2, we explored the Lorenz ratio ($\frac{L}{L_{0}}$) by using non-fluid (NF) type expressions. Fig. 1(a) and (b) have shown the Lorenz ratio versus $\frac{\epsilon_{F}}{T}$ for 3D-G and 2D-G systems respectively. By using the non-fluid (NF) type of Eq. (94) and its Sommerfeld limit Eq. (98), we plotted the black solid and dotted lines in Fig. 1(a). We used the specific heat ${[C_{V}]_{e1}}=\frac{\partial}{\partial T}\Big(\frac{U}{N}\Big)$, which is traditionally used in solid state or non-fluid systems. We are getting $L/L_{0}=1$ in the limit of $\epsilon_{F}/T\gg 1$ or Fermi Liquid (FL) regime of graphene case, which is as expected for most of the metal cases. The deviation of the Lorenz ratio from 1 has been started below $\frac{\epsilon_{F}}{T}\approx 10$ as we notice from the black solid line. We have also plotted the red solid line by using Eq. (95), where another definition of specific heat $C_{V}]_{e2}$} is used in the thermal conductivity expression. In the case of SL of 3D-G (the red dotted line), the Lorenz ratio is three instead of one. For the general case (the red solid line), the Lorenz ratio deviates from this constant line $3$ below $\frac{\epsilon_{F}}{T}\approx 10$. For the 2D-G system, similar kinds of black and red solid lines are plotted in Fig. 1(b) by using Eq. (12) and Eq. (13). Their respective SLs by using Eq. (14) and Eq. (15), are plotted as black and red horizontal dotted lines. We considered the surface area $(S)$ in place of volume $(V)$ in the 2D system. The Lorenz ratio when we take $C_{V}]_{e2}$}, the ratio becomes 2 (the red-dotted line).
@@
The Fig. 2(a) and (b), similar to Fig. 1, display the Lorenz ratio in terms of $\frac{\epsilon_{F}}{T}$ for non-fluid expressions of NR case. Though the real metals have a constant chemical potential within the range, $\epsilon_{F}=2-10$ eV, we imagine that it can be tunable from 0 to $\infty$. In Fig. 2(a), we notice that the WF law is valid in the SL of the 3D metal case, expressed by Eq. (84) (green dot-dashed line). We have taken $T=60^{\circ}$ K ($\approx 0.005$ eV), so metal range $\epsilon_{F}=2$-$10$ eV will give the ratio $\frac{\epsilon_{F}}{T}=400$-$2000$ ($\gg 1$), which is quite good domain for SL. Therefore, being situated in the SL domain, metal always follows the WF law. A slight larger value $L=1.5L_{0}$ comes from Eq. (88) (blue dotted line), when specific heat $C_{V}]_{e2}$} is taken. A similar kind of behavior can be seen in Fig. 2(b) for 2D-NR systems, but interestingly, SLs in both cases, given by Eqs. (105) and (106) are coincided at $L=L_{0}$ (blue-dotted and green dot-dashed lines). It probably indicates that if one builds 2D metal systems, then it will also follow the WF law as their Fermi energy range is again expected to be situated in the SL domain.
@@
Before going to the next graphical results, let us briefly analyze what we have learned from Fig. 1(a), (b) and Fig. 2(a), (b). Grossly, the non-fluid type expressions of Lorenz ratios for all cases - 3D-NR, 2D-NR, 3D-G, and 2D-G are saturated near 1 in the SL range, which means that the WF law is obeyed in the SL range. Among all four cases, 3D-NR and 2D-G are realistic cases. The former is applicable to metal systems, whose Fermi energies remain constant and within the SL range, so they always follow the WF law. Latter case - the non-fluid expressions of Lorenz ratio for 2D-G can be applicable for realistic graphene system with high doping concentration ($\epsilon_{F}$ is quite large or $\epsilon_{F}/T\gg 1$), where the WF law is obeyed well again [crossno2016observation]. However, by decreasing the doping concentration, when the graphene system approaches the charge-neutral points ($\epsilon_{F}$ is quite small or $\epsilon_{F}/T\ll 1$), the WF law violation domain is observed [crossno2016observation]. According to our non-fluid expressions, the WF law violation is also expected in the $\epsilon_{F}/T\ll 1$ domain but the outcome $L/L_{0}<1$ is not matching with experimental outcome $L/L_{0}>1$ by Crossno et al. [crossno2016observation]. A possible explanation of $L/L_{0}>1$ in the appearance of electron hydrodynamics [Theory_G_WF] property in the domain $\epsilon_{F}/T<1$, which is popularly called the Dirac fluid domain. Present work is firstly intended to highlight the overlooked possibility of Wiedemann-Franz law violation of any systems like NR, G or UR even we consider the non-fluid expression of $\kappa$ and $\sigma$. Then, its next attempt is to show the transition from non-fluid (NF) to fluid (F) pictures in a systematic way. An approach towards the fluid description is addressed in formalism part Sec: (II.2). Let us come to the numerical estimation for fluid-type formalism.
@@
For numerical curves of fluid-type expressions, let us focus only on the 2D-G case. Although formalism of other cases with an Appendix is provided in the present article, one may go for result generation for graphene case. Using Eqs. (30) and (31), black and red solid lines are plotted in the left panel of Fig. 3. They are saturated at the values of 2 and 4, respectively, in the high doping or $\epsilon_{F}/T\gg 1$ domain. It means that if the fluid type expression is valid in this high doping domain 2D-G system, then the WF law may not be obeyed. However, the experimental data [crossno2016observation] claims that the WF law is well obeyed in the high doping or $\epsilon_{F}/T\gg 1$ domain. This means that the non-fluid expression of the Lorenz ratio should be applicable to this domain instead of fluid-type expression. The SL of Eqs. (30) and (31) will also be
@@
which are plotted by black and red dotted lines, respectively, in the left panel of Fig. 3. They blow up from their saturated values (2 and 4) when $\epsilon_{F}/T$ or doping concentration decreases. The experimental data of Crossno et al. [crossno2016observation] is pasted, where S1 and S2 represent the samples of the 2D-G system in terms of lower to higher values of minimum charge carrier density. Theoretically, zero charge carrier density can be reached at $\epsilon_{F}/T\rightarrow 0$ but experimentally, a minimum charge carrier density [crossno2016observation] will be achieved. So, from Fig.3(a), we notice that our fluid-based expression of $\frac{L_{G2}^{2D}}{L_{0}}$, given in Eq.(30), (31), can’t be able to explain the experimental data points. It indicates that our formalism needs next level modification. Due to this step by step development of our formalism from (standard) non-fluid to (new directional) fluid description, we put the name of Sec: (II.2) as ”towards the fluid description”. Now if we revisit our RTA based kinetic theory formalism, then we can find that we have systematically considered the local thermalization concept of fluid. In this concept, we consider fluid as a large number fluid elements, having velocity $u^{\mu}$, temperature $T$ and chemical potential $\epsilon_{F}$. Now, fluid elements will be at different position and time $(x^{\mu})$, so, considering $u^{\mu}(x^{\mu})$, $T(x^{\mu})$, $\epsilon_{F}(x^{\mu})$ as function of $x^{\mu}$ can map mathematically the local thermalization concept. Since present work is only focus on thermal and electrical conductivity, so only $T(x^{\mu})$ is considered and others $u^{\mu}(x^{\mu})$ and $\epsilon_{F}(x^{\mu})$ are ignored. Reader can see this consideration in Eq.(20), where the information of temperature gradient $\frac{\partial T}{\partial x^{i}}$ is entered. Now the another term specific heat $C_{V}]_{e}$ $\approx\frac{\partial u}{\partial T}$, which is basically considered as temperature derivative of average energy of electron. This average energy of electron information can be modified into enthalpy per electron $h=\frac{u+P}{n}$ in fluid picture. Another point is that electron-hole plasma picture will be important when we go $\frac{\epsilon_{F}}{T}<1$ domain (towards charge neutrality point). So, we should take enthalpy per net electron $h=\frac{u+P}{n}$, where $u=u_{e}+u_{h}$, $P=P_{e}+P_{h}$ and $n=n_{e}-n_{h}$ will be respectively total internal energy density, pressure and net charge density. We have adopted the replacement
$C_{V}]_{e}\approx\frac{\partial u}{\partial T}\,\rightarrow\,\frac{h}{T}=\frac{u+P}{nT}$ (34)
@@
$\left(\frac{L_{G1}^{2D}}{L_{0}}\right)=\frac{6}{\pi^{2}}\left[C_{V}\right]^{F/NF}_{G}\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}~.$ (36)
@@
  • Fluid (F) domain: Fluid based Lorenz ratio, given in Eq.(36) valid well with experimental data in the range $0.2<\frac{\epsilon_{F}}{T}<0.7$.
  • Non-Fluid (NF) domain: Non-Fluid based Lorenz ratio, given in Eq.(12), valid well with experimental data in the ranges (a) $2.5<\frac{\epsilon_{F}}{T}<10$ and (b) $10<\frac{\epsilon_{F}}{T}<\infty$. We can called subrange (b) as ohmic non-fluid or NF(Ohmic) domain, since we get WF law or $\frac{L}{L_{0}}=1$ for this $\frac{\epsilon_{F}}{T}>10$ domain, where Fermi-Dirac distribution function can be assumed as step function. However, for subrange (a), we get $\frac{L}{L_{0}}<1$ because assumption of step function is not further valid here. Corresponding domains are also marked in Fig.3(b), where we can notice that $C_{V}^{e}\propto\frac{T}{\epsilon_{F}}$ follows in NF(Ohmic) domain (due to validation of electron step function like distribution) but a transition from $C_{V}^{e}\propto\frac{T}{\epsilon_{F}}$ to $C_{V}^{e}=2k_{B}\Bigg[3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg]$ is observed in NF domain ($2.5<\frac{\epsilon_{F}}{T}<10$). Reader can understand this transition from Eq.(10).
  • Mixed or (F+NF) domain: The domain $0.7<\frac{\epsilon_{F}}{T}<2.5$ is marked as mixed or (F+NF) domain as we have to go with replacement $C_{V}^{e}\rightarrow\frac{u+P}{nT}$ or transit from NF based quantity $C_{V}^{e}$ to fluid-based quantity $\frac{u+P}{nT}$. This domain may be considered as a grey zone and we may put an open question for further research- How this transition happens? Interestingly, reader can notice a $\frac{\epsilon_{F}}{T}$ independent trend approximately of $C_{V}^{e}$ and $\frac{u+P}{nT}$ within this (F+NF) domain.
@@
$\left(\frac{L}{L_{0}}\right)_{F/NF}^{G}=\frac{6}{\pi^{2}}\,\big[C_{V}\big]_{F/NF}\,\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\,(A=e^{x})~,$ (40)
$\left(\frac{L}{L_{0}}\right)_{F/NF}^{Q}=\frac{9}{\pi^{2}}\,\big[C_{V}\big]_{F/NF}\,\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}~,$ (41)
@@
Using Eq.(41), we have plotted blue and red solid curves to fit experimental data point $S_{1}$(blue circles) and $S_{2}$(red circles). For blue line, tuning parameters are taken as $x_{F}=0.2$, $x_{NF}=0.8$, $b_{F}=2$, $b_{NF}=0.7$ and for red line, tuning parameters are taken as $x_{F}=0.9$, $x_{NF}=2$, $b_{F}=2$, $b_{NF}=0.7$. By using a better version of x-dependent switching functions, one may go with better fitting of experimental data.
Figure 5: Lorenz ratio of 2D-G system versus number density for different theoretical works by Rycerz [rycerz2021wiedemann], Tu et al. [tu_Yin_23_the], Lucus et al. [Theory_G_WF], including present work
However, instead of quantitative exact fitting, the core aim of the present work may be considered as its qualitative message or understanding on the F and NF domain and their transition.
@@
This qualitative message of transition from NF to F is in agreement with earlier theoretical works by Lucas et al. [Theory_G_WF]. Though final expressions of Lorenz ratio of Ref. [Theory_G_WF] and the present work are little different but both works have similar qualitative message- NF to F transition is associated with the WF law violation and enhancement of Lorenz ratio more than one. There are other works also which provide little different reasons for getting WF law violations. Ref. [tu_Yin_23_the] has linked this fact with presence of band gap and bipolar diffusion, while Ref. [rycerz2021wiedemann] has built thermodynamical expression of electron-hole plasma in terms of polylogarithm function, which can provide a greater than one values of Lorenz ratio. All of these theoretical results for $\frac{L}{L_{0}}$ along with present work are compiled in Fig.5 for a comparative visibility.
[…]
Figure 6: Our present theoretical results of electrical conductivity (LEFT) and thermal conductivity (RIGHT) with different theoretical predictions [Theory_G_WF, PhysRevB.76.144502] accompanied by experimental data  [crossno2016observation] versus number density
IV Brief Interpretation of Results